The average distance between Earth and the Sun is 2.5 × 1011m. (a) Calculate the average speed of Earth in its orbit (assumed to be circular) in meters per second. (b) What is this speed in miles per hour?


1
Expert's answer
2021-06-15T15:37:46-0400

a)


"v=\frac{2\pi r}{T}\\v=\frac{2\pi (2.5\cdot10^{11})}{365.2\cdot24\cdot3600}=5.0\cdot10^{4}\frac{m}{s}"

b)


"v=1.1\cdot10^{5}\ mph"